# Constraints on Long-Ranged Interactions Between Dark Matter and the Standard Model

### v1

@@ I Introduction
New interactions between dark matter and the Standard Model can take a variety of forms [1], with different dark matter models and forces constrained by different experimental and astrophysical observations. At low energies, however, most such interactions reduce to one of a few possible classical potentials [2], with the nature of the mediator determining parameters such as the potential’s range and species-dependent couplings.
@@
Such long-range forces can have major implications for our understanding of astrophysics. In fact, even short-range interactions of dark matter particles can have major implications for dark matter structures, as studies of self-interacting dark matter have shown [4, 5, 6, 7]. Long-range forces, however, can have much larger effects, thanks to the potential for collective effects of many dark matter particles causing or experiencing a force coherently. As a result, long-range forces—both between dark matter particles alone, and between dark matter particles and the Standard Model—can have far more dramatic consequences for astrophysical structures. In particular, our understanding of the distribution of dark matter throughout the universe—whether in the CMB or within galaxies and galaxy cluster—is based on its interactions with the Standard Model, assuming that the only such interaction is gravity. Additional long-range interactions of dark matter and the Standard Model could therefore dramatically modify the dark matter densities that we infer. In this work, we study both this effect and other implications of such a new force.
@@
$V$ $=q_{\rm S}q_{\rm D}\alpha_{\rm SD}G\frac{m_{\rm S}m_{\rm D}}{r}e^{-r/\lambda}$ (1)
where $\alpha_{\rm SD}$ is the characteristic strength of the new force compared to gravity, $G$ is Newton’s constant, $m_{\rm S}$ ($m_{\rm D}$) is the mass of the Standard Model (dark matter) particle, and $q_{\rm S}$ ($q_{\rm D}$) is the charge of that particle, which we will assume to be order-one but potentially with either sign.
Crucially, a new Standard Model-dark matter force, as in Eq. (1), involves a mediator coupled to both Standard Model particles and dark matter particles, and therefore also leads to a new force between Standard Model particles alone and between dark matter particles alone. These forces need not have similar strengths: letting $g_{S}$ ($g_{D}$) be the coupling to a Standard Model (dark matter) particle of charge $q_{S}=1$ ($q_{D}=1$), we have two independent degrees of freedom setting the three characteristic strengths: $\alpha_{\rm SS}=g_{S}^{2}/(4\pi)$, $\alpha_{\rm DD}=g_{D}^{2}/(4\pi)$, and $\alpha_{\rm SD}=g_{S}g_{D}/(4\pi)$, with $\alpha_{\rm SS}$ ($\alpha_{\rm DD}$) defined analogously to $\alpha_{\rm SD}$ but for Standard Model-Standard Model (dark matter-dark matter) forces
@@
The remainder of this work is organized as follows: In Section II, we present constraints on new long-ranged interactions between dark matter and the Standard Model from combining constraints on dark matter-only and Standard Model-only forces, since any force between the two must necessarily couple to both. We therefore briefly review these individual constraints and present the resulting constraints on $\alpha_{\rm SD}$. We then describe a new constraint directly on $\alpha_{\rm SD}$ based on dynamical friction in UFDs in Section III. We begin by summarizing the observation status of UFDs and the effects of dynamical friction itself on the evolution of stellar distributions within them. We then consider a couple of related subtleties, including the modification of UFD dark matter distributions by the presence of a new force and the influence of plasma effects on our results. Finally, in Section IV, we summarize the observational status of long-ranged forces between dark matter and the Standard Model.
II Combined Constraints
@@
$\alpha_{\rm SD}=\sqrt{\alpha_{\rm SS}\alpha_{\rm DD}};$ (2)
see the discussion below Eq. (1). We will refer to these as “combined constraints” for convenience. For many classes of new forces, these will set the strongest constraints in this work, thanks in large part to the extremely strong limits which can be placed on new long-range interactions of Standard Model particles. We therefore begin by reviewing the present state of constraints on $\alpha_{\rm SS}$ and $\alpha_{\rm DD}$, before turning to limits set directly on dark matter-Standard Model interactions in the next section.
@@
Finally, we note that—in the case of forces where both signs of charge are present within dark matter, such that the dark matter forms a plasma—one additional effect may lead to significantly stronger constraints in the foreseeable future: plasma instabilities. In particular, Ref. [11] shows that collisional instabilities would lead to exponential growth of perturbations in various astrophysical systems for many new force parameters of interest. While this has not been used to set constraints to-date, it may be possible to do so in the future; see Section III.4 for further discussion.
New long-ranged forces acting on the Standard Model alone have been constrained with a variety of experiments and observations, although their interpretation is complicated by the question of how similar they are to gravity. In particular, long-ranged forces are generally easier to constrain when they lead to larger apparent violations of the equivalence principle. This is because equivalence principle-violating effects can be detected via differential accelerations between different materials, whereas equivalence principle-preserving ones can be detected only through the difference between the $r^{-1}$ behavior of gravity and the Yukawa potential of Eq. (1), which is only weakly suppressed on length scales $r<\lambda$.
@@
A variety of models with new gravity-like forces have been proposed, including ones based on scalar gravity, bigravity, dilatons, etc. We will not attempt to review these models, and the extent to which the resulting forces can be equivalence principle-conserving, in this work; see, for example, Refs. [12, 13, 14, 15]. We will instead present constraints based on the two limiting cases of a new force’s equivalence principle violation: order-one violation (i.e., in the language of Eq. (1), a force which couples with $q_{S}\sim 1$ for one test mass but $q_{S}\ll 1$ for the other) and zero violation (i.e. a force exactly proportional to any Standard Model test object’s total mass, with $q_{S}=1$ always). Typical models of new forces will then lie somewhere between these extremes, and the resulting constraints can generally be obtained straightforwardly by interpolating between these limits.
@@
$\left|\alpha_{\rm SS}^{\rm(max.~{}EPV)}\right|\lesssim{\rm few}\times 10^{-15},$ (3)
independent of the new force’s range for the ranges we consider, which are all much larger than the cylinders’ distance from Earth’s surface ($\sim 700$ km). As we noted above, many models of new forces lead to equivalence principle violation far below maximal; in this case, the limit of Eq. (3) should be weakened in proportion to the fractional acceleration difference expected between titanium and platinum for that model.
@@
$\left|\alpha_{\rm SS}^{\rm(no~{}EPV)}\right|\leq 1\qquad(\lambda\gg 100~{}\mu{\rm m}).$ (4)
@@
$\begin{split}V_{\rm new}&=\alpha_{\rm SS}G\frac{m_{1}m_{2}}{r}e^{-r/\lambda}\\ &\approx\alpha_{\rm SS}Gm_{1}m_{2}\left(\frac{1}{r}-\frac{1}{\lambda}+\frac{r}{2\lambda^{2}}+...\right).\end{split}$ (5)
@@
$\left|\alpha_{\rm SS}^{\rm(no~{}EPV)}\right|\lesssim 7\times 10^{-10}\left(\frac{\lambda}{\rm AU}\right)^{2}\qquad(\lambda\gg{\rm AU}).$ (6)
@@
Figure 1: Existing constraints on new long-range forces on Standard Model particles (top) and on dark matter particles (bottom), i.e. on $\alpha_{\rm SS}$ and on $\alpha_{\rm DD}$; see Section II for discussion. Top: Shown are constraints on SM-SM forces in both the maximal and zero equivalence principle violation limits. In the presence of non-negligible equivalence principle violation, the strongest constraint at all ranges comes from the MICROSCOPE experiment [16]. In the absence of any EP violation, the constraints arise from planetary observations [18, 19, 20] or, once the range is too large to distinguish the force from gravity on observed length scales, from the known strength of gravity. Bottom: A summary of constraints on a new, long-range, DM-DM force. One constraint shown arises from incoherent scattering in the Bullet Cluster [21, 22, 23, 24] (where we have shown two example dark matter masses) which is applicable for attractive, repulsive, and net-neutral dark matter. Also shown is a constraint from coherent interactions in the bullet cluster which is applicable for attractive forces [9]. An additional constraint shown is from effects of attractive forces on the formation of large-scale structure [10]. A final existing constraint is from self-binding in the presence of UFDs and applies to repulsive forces only [9]. Also shown is the potential constraint that could be obtained in the future from excluding exponentially growing plasma instabilities in astrophysical systems, in the case of net-neutral dark matter [11].
III Dynamical Friction in Ultrafaint Dwarf Galaxies
@@
$\alpha_{\rm SD}^{\rm(max.~{}EPV)}\lesssim 10^{-4}\qquad(\lambda\gg 8~{}{\rm kpc})$ (7)
with $8$ kpc the distance to the Milky Way center. This can likely be extended to shorter ranges, albeit suppressed by a factor of $(8~{}{\rm kpc}/\lambda)^{2}$, based on the effects of dark matter gradients; see Ref. [9]. We do not consider this in detail, however, since this constraint is subleading compared to the combined constraints discussed in Section II. Some other constraints on $\alpha_{\rm SD}$ are presented in Ref. [8], but we will not review them here since they are generally subdominant and carry some additional model dependence.
III.1 Properties of Ultrafaint Dwarfs
@@
The UFD that we will focus on here is Segue 1, whose properties are comparatively well-known [27, 28, 29, 30]. We note, however, that our analysis can be straightforwardly extended to many other UFDs as well. Segue 1 is a UFD located approximately $23$ kpc from the Sun, with a luminosity of roughly $300~{}L_{\odot}$ (and thus $\sim 300$ stars), a half-light radius of approximately $32.3$ pc, and a stellar line-of-sight velocity dispersion of $3.7$ km s-1. Simulations of UFD population evolution predict that Segue I likely has a mass of approximately $10^{9}~{}M_{\odot}$, while estimates of its age, based largely on the metallicity of its stars, suggest an age of order 10 Gyr [31, 32, 33]. We will adopt these values for our calculations (up to corrections discussed in Section III.3), although we note that their uncertainties vary from several percent to order-one. None of these uncertainties should affect our conclusions by more than order-one factors, however.
@@
$\rho_{\chi}(r)=\frac{(3-\gamma)M_{\chi,{\rm tot}}}{4\pi R_{\chi}^{3}}{\left(\frac{r}{R_{\chi}}\right)}^{-\gamma}\left(1+\frac{r}{R_{\chi}}\right)^{\gamma-4}$ (8)
@@
$\rho_{\chi}(r)=\frac{M_{\chi,{\rm tot}}}{4\pi R_{\chi}^{3}\mathcal{N}}{\left(\frac{r}{R_{\chi}}\right)}^{-1}\left(1+\frac{r}{R_{\chi}}\right)^{-2}$ (9)
@@
$\rho_{*}(r)=\frac{3M_{*,{\rm tot}}}{4\pi R_{*}^{3}}\left(1+\frac{r^{2}}{R_{*}^{2}}\right)^{-5/2},$ (10)
@@
In this section, we consider the effects of gravity alone, without any new forces; we account for the effects of the new forces of interest in this work in Section III.3.
@@
$\left\langle\sigma_{*,{\rm los}}^{2}\right\rangle\approx\frac{GM_{\frac{1}{2}}}{3R_{\frac{1}{2}}}$ (11)
@@
$\frac{dP_{*}}{dr}=\frac{d}{dr}{\left(\frac{T_{*}(r)\rho_{*}(r)}{m_{*}}\right)}=-\frac{G\rho_{*}(r)}{r^{2}}M_{\rm tot}(r)$ (12)
@@
$\frac{dP_{\chi}}{dr}=\frac{d}{dr}{\left(\frac{T_{\chi}(r)\rho_{\chi}(r)}{m_{*}}\right)}=-\frac{G\rho_{\chi}(r)}{r^{2}}M_{\rm tot}(r)$ (13)
@@
$\sigma_{*}(R_{*})$ $=\sqrt{\frac{2GM_{\chi,{\rm tot}}R_{*}^{2}}{3R_{\chi}^{3}}}$ (14a)
$\sigma_{\chi}(R_{*})$ $=\sqrt{\frac{GM_{\chi,{\rm tot}}}{30R_{\chi}}}$ (14b)
@@
$U_{**}$ $=\frac{3\pi GM_{*,{\rm tot}}^{2}}{32R_{*}}$ (15a)
$U_{*\chi}$ $=\frac{4GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{\pi^{3/2}R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&\frac{3}{2}&2\\[2.0pt] \frac{5}{2}&\frac{5}{2}&3\end{matrix}\right.\right)$ (15b)
$U_{\chi*}$ $=\frac{4GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{\pi^{3/2}R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&1&\frac{3}{2}\\[2.0pt] \frac{3}{2}&\frac{5}{2}&3\end{matrix}\right.\right)$ (15c)
@@
III.2 Effects of Dynamical Friction

### v2

@@ 1 Introduction
New interactions between dark matter and the Standard Model can take a variety of forms [1], with different dark matter models and forces constrained by different experimental and astrophysical observations. At low energies, however, most such interactions reduce to one of a few possible classical potentials [2] (see e.g. Ref. [3] for an exception), with the nature of the mediator determining parameters such as the potential’s range and species-dependent couplings.
@@
Such long-range forces can have major implications for our understanding of astrophysics. In fact, even short-range interactions of dark matter particles can have major implications for dark matter structures, as studies of self-interacting dark matter have shown [5, 6, 7, 8]. Long-range forces, however, can have much larger effects, thanks to the potential for collective effects of many dark matter particles causing or experiencing a force coherently (see, for example, Ref. [9]). As a result, long-range forces—both between dark matter particles alone, and between dark matter particles and the Standard Model—can have far more dramatic consequences for astrophysical structures. In particular, our understanding of the distribution of dark matter throughout the universe—whether in the CMB or within galaxies and galaxy cluster—is based on its interactions with the Standard Model, assuming that the only such interaction is gravity. Additional long-range interactions of dark matter and the Standard Model could therefore dramatically modify the dark matter densities that we infer. In this work, we study both this effect and other implications of such a new force.
@@
$V$ $=q_{\rm S}q_{\rm D}\alpha_{\rm SD}G\frac{m_{\rm S}m_{\rm D}}{r}e^{-r/\lambda}$ (1.1)
where $\alpha_{\rm SD}$ is the characteristic strength of the new force compared to gravity, $G$ is Newton’s constant, $m_{\rm S}$ ($m_{\rm D}$) is the mass of the Standard Model (dark matter) particle, and $q_{\rm S}$ ($q_{\rm D}$) is the charge of that particle, which we will assume to be order-one but potentially with either sign. Note that our parametrization of the strength relative to gravity means that we will frequently consider $\alpha_{\rm SD}\gg 1$ in this work. Such couplings will nonetheless generally be perturbative; we discuss exceptions in Appendix B.
Crucially, a new Standard Model-dark matter force, as in Eq. (1.1), involves a mediator coupled to both Standard Model particles and dark matter particles, and therefore also leads to a new force between Standard Model particles alone and between dark matter particles alone. These forces need not have similar strengths: letting $g_{S}$ ($g_{D}$) be the coupling to a Standard Model (dark matter) particle of charge $q_{S}=1$ ($q_{D}=1$), we have two independent degrees of freedom setting the three characteristic strengths: $\alpha_{\rm SS}=g_{S}^{2}/(4\pi)$, $\alpha_{\rm DD}=g_{D}^{2}/(4\pi)$, and $\alpha_{\rm SD}=g_{S}g_{D}/(4\pi)$, with $\alpha_{\rm SS}$ ($\alpha_{\rm DD}$) defined analogously to $\alpha_{\rm SD}$ but for Standard Model-Standard Model (dark matter-dark matter) forces
@@
The remainder of this work is organized as follows: In Section 2, we present constraints on new long-ranged interactions between dark matter and the Standard Model from combining constraints on dark matter-only and Standard Model-only forces, since any force between the two must necessarily couple to both. We therefore briefly review these individual constraints and present the resulting constraints on $\alpha_{\rm SD}$. We then describe a new constraint directly on $\alpha_{\rm SD}$ based on dynamical friction in UFDs in Section 3. We begin by summarizing the observation status of UFDs and the effects of dynamical friction itself on the evolution of stellar distributions within them. We then consider a couple of related subtleties, including the modification of UFD dark matter distributions by the presence of a new force and the influence of plasma effects on our results. Finally, in Section 4, we summarize the observational status of long-ranged forces between dark matter and the Standard Model.
2 Combined Constraints
@@
$\alpha_{\rm SD}=\sqrt{\alpha_{\rm SS}\alpha_{\rm DD}};$ (2.1)
see the discussion below Eq. (1.1). We will refer to these as “combined constraints” for convenience. For many classes of new forces, these will set the strongest constraints in this work, thanks in large part to the extremely strong limits which can be placed on new long-range interactions of Standard Model particles. We therefore begin by reviewing the present state of constraints on $\alpha_{\rm SS}$ and $\alpha_{\rm DD}$, before turning to limits set directly on dark matter-Standard Model interactions in the next section.
@@
Finally, we note that—in the case of forces where both signs of charge are present within dark matter, such that the dark matter forms a plasma—one additional effect may lead to significantly stronger constraints in the foreseeable future: plasma instabilities. In particular, Ref. [13] shows that collisional instabilities would lead to exponential growth of perturbations in various astrophysical systems for many new force parameters of interest. While this has not been used to set constraints to-date, it may be possible to do so in the future; see Section 3.4 for further discussion.
New long-ranged forces acting on the Standard Model alone have been constrained with a variety of experiments and observations, although their interpretation is complicated by the question of how similar they are to gravity. In particular, long-ranged forces are generally easier to constrain when they lead to larger apparent violations of the equivalence principle. This is because equivalence principle-violating effects can be detected via differential accelerations between different materials, whereas equivalence principle-preserving ones can be detected only through the difference between the $r^{-1}$ behavior of gravity and the Yukawa potential of Eq. (1.1), which is only weakly suppressed on length scales $r<\lambda$.
@@
A variety of models with new gravity-like forces have been proposed, including ones based on scalar gravity, bigravity, dilatons, etc. We will not attempt to review these models, and the extent to which the resulting forces can be equivalence principle-conserving, in this work; see, for example, Refs. [14, 15, 16, 17]. We will instead present constraints based on the two limiting cases of a new force’s equivalence principle violation: order-one violation (i.e., in the language of Eq. (1.1), a force which couples with $q_{S}\sim 1$ for one test mass but $q_{S}\ll 1$ for the other) and zero violation (i.e. a force exactly proportional to any Standard Model test object’s total mass, with $q_{S}=1$ always). Typical models of new forces will then lie somewhere between these extremes, and the resulting constraints can generally be obtained straightforwardly by interpolating between these limits.
@@
$\left|\alpha_{\rm SS}^{\rm(max.~{}EPV)}\right|\lesssim{\rm few}\times 10^{-15},$ (2.2)
independent of the new force’s range for the ranges we consider, which are all much larger than the cylinders’ distance from Earth’s surface ($\sim 700$ km). As we noted above, many models of new forces lead to equivalence principle violation far below maximal; in this case, the limit of Eq. (2.2) should be weakened in proportion to the fractional acceleration difference expected between titanium and platinum for that model.
@@
$\left|\alpha_{\rm SS}^{\rm(no~{}EPV)}\right|\leq 1\qquad(\lambda\gg 100~{}\mu{\rm m}).$ (2.3)
@@
$\begin{split}V_{\rm new}&=\alpha_{\rm SS}G\frac{m_{1}m_{2}}{r}e^{-r/\lambda}\\ &\approx\alpha_{\rm SS}Gm_{1}m_{2}\left(\frac{1}{r}-\frac{1}{\lambda}+\frac{r}{2\lambda^{2}}+...\right).\end{split}$ (2.4)
@@
$\left|\alpha_{\rm SS}^{\rm(no~{}EPV)}\right|\lesssim 7\times 10^{-10}\left(\frac{\lambda}{\rm AU}\right)^{2}\qquad(\lambda\gg{\rm AU}).$ (2.5)
@@
Figure 1: Existing constraints on new long-range forces on Standard Model particles (top) and on dark matter particles (bottom), i.e. on $\alpha_{\rm SS}$ and on $\alpha_{\rm DD}$; see Section 2 for discussion. Top: Shown are constraints on SM-SM forces in both the maximal and zero equivalence principle violation limits. In the presence of non-negligible equivalence principle violation, the strongest constraint at all ranges comes from the MICROSCOPE experiment [18]. In the absence of any EP violation, the constraints arise from planetary observations [20, 21, 22] or, once the range is too large to distinguish the force from gravity on observed length scales, from the known strength of gravity. Bottom: A summary of constraints on a new, long-range, DM-DM force. One constraint shown arises from incoherent scattering in the Bullet Cluster [23, 24, 25, 26] (where we have shown two example dark matter masses) which is applicable for attractive, repulsive, and net-neutral dark matter. Also shown is a constraint from coherent interactions in the bullet cluster which is applicable for attractive forces [11]. An additional constraint shown is from effects of attractive forces on the formation of large-scale structure [12]. A final existing constraint is from self-binding in the presence of UFDs and applies to repulsive forces only [11]. Also shown is the potential constraint that could be obtained in the future from excluding exponentially growing plasma instabilities in astrophysical systems, in the case of net-neutral dark matter [13].
3 Dynamical Friction in Ultrafaint Dwarf Galaxies
@@
$\alpha_{\rm SD}^{\rm(max.~{}EPV)}\lesssim 10^{-4}\qquad(\lambda\gg 8~{}{\rm kpc})$ (3.1)
with $8$ kpc the distance to the Milky Way center. This can likely be extended to shorter ranges, albeit suppressed by a factor of $(8~{}{\rm kpc}/\lambda)^{2}$, based on the effects of dark matter gradients; see Ref. [11]. We do not consider this in detail, however, since this constraint is subleading compared to the combined constraints discussed in Section 2. Some other constraints on $\alpha_{\rm SD}$ are presented in Ref. [10, 28], but we will not review them here since they are generally subdominant and carry some additional model dependence.
3.1 Properties of Ultrafaint Dwarfs
@@
The UFD that we will focus on here is Segue 1, whose properties are comparatively well-known [30, 31, 32, 33]. We note, however, that our analysis can be straightforwardly extended to many other UFDs as well. Segue 1 is a UFD located approximately $23$ kpc from the Sun, with a luminosity of roughly $300~{}L_{\odot}$ (and thus $\sim 300$ stars), a half-light radius of approximately $32.3$ pc, and a stellar line-of-sight velocity dispersion of $3.7$ km s-1. Simulations of UFD population evolution predict that Segue I likely has a mass of approximately $10^{9}~{}M_{\odot}$, while estimates of its age, based largely on the metallicity of its stars, suggest an age of order 10 Gyr [34, 35, 36]. We will adopt these values for our calculations (up to corrections discussed in Section 3.3), although we note that their uncertainties vary from several percent to order-one. None of these uncertainties should affect our conclusions by more than order-one factors, however.
@@
$\rho_{\chi}(r)=\frac{(3-\gamma)M_{\chi,{\rm tot}}}{4\pi R_{\chi}^{3}}{\left(\frac{r}{R_{\chi}}\right)}^{-\gamma}\left(1+\frac{r}{R_{\chi}}\right)^{\gamma-4}$ (3.2)
@@
$\rho_{\chi}(r)=\frac{M_{\chi,{\rm tot}}}{4\pi R_{\chi}^{3}\mathcal{N}}{\left(\frac{r}{R_{\chi}}\right)}^{-1}\left(1+\frac{r}{R_{\chi}}\right)^{-2}$ (3.3)
@@
$\rho_{*}(r)=\frac{3M_{*,{\rm tot}}}{4\pi R_{*}^{3}}\left(1+\frac{r^{2}}{R_{*}^{2}}\right)^{-5/2},$ (3.4)
@@
These profiles should be truncated at some radius due to the effects of tidal forces [41]. While the tidal radius of Segue I is not known precisely, it is expected to be at least of order $R_{\chi}$, given the fairly small eccentricity of Segue I’s orbit [35]. This truncation is therefore negligible for our purposes, since it affects only dark matter much farther from the stars than the largest impact parameters we will consider (see Section 3.2 and Appendix B).
In this section, we consider the effects of gravity alone, without any new forces; we account for the effects of the new forces of interest in this work in Section 3.3.
@@
$\left\langle\sigma_{*,{\rm los}}^{2}\right\rangle\approx\frac{GM_{\frac{1}{2}}}{3R_{\frac{1}{2}}}$ (3.5)
@@
$\frac{dP_{*}}{dr}=\frac{d}{dr}{\left(\frac{T_{*}(r)\rho_{*}(r)}{m_{*}}\right)}=-\frac{G\rho_{*}(r)}{r^{2}}M_{\rm tot}(r)$ (3.6)
@@
$\frac{dP_{\chi}}{dr}=\frac{d}{dr}{\left(\frac{T_{\chi}(r)\rho_{\chi}(r)}{m_{*}}\right)}=-\frac{G\rho_{\chi}(r)}{r^{2}}M_{\rm tot}(r)$ (3.7)
@@
$\sigma_{*}(R_{*})$ $=\sqrt{\frac{2GM_{\chi,{\rm tot}}R_{*}^{2}}{3R_{\chi}^{3}}}$ (3.8a)
$\sigma_{\chi}(R_{*})$ $=\sqrt{\frac{GM_{\chi,{\rm tot}}}{30R_{\chi}}}$ (3.8b)
@@
$U_{**}$ $=\frac{3\pi GM_{*,{\rm tot}}^{2}}{32R_{*}}$ (3.9a)
$U_{*\chi}$ $=\frac{4GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{\pi^{3/2}R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&\frac{3}{2}&2\\[2.0pt] \frac{5}{2}&\frac{5}{2}&3\end{matrix}\right.\right)$ (3.9b)
$U_{\chi*}$ $=\frac{4GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{\pi^{3/2}R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&1&\frac{3}{2}\\[2.0pt] \frac{3}{2}&\frac{5}{2}&3\end{matrix}\right.\right)$ (3.9c)
@@
3.2 Effects of Dynamical Friction